OpenStudy (anonymous):

The vectors a=2i−3j+k,b=4i+3j−2k and c=2i−4j+xk are linearly dependent. Find the value of x.

OpenStudy (anonymous):

\[\vec{a}=2\vec{i}-3\vec{j}+\vec{k}\\ \vec{b}=4\vec{i}+3\vec{j}-2\vec{k}\\ \vec{c}=2\vec{i}-4\vec{j}+x\vec{k}\] For these vectors to be linearly dependent, there must be some non-zero constant n_i (i = 1,2,3) such that the following is satisfied: \[n_1\vec{a} + n_2\vec{b}+n_3\vec{c}=\vec{0}\\ \small n_1\left(2\vec{i}-3\vec{j}+\vec{k}\right) + n_2\left(4\vec{i}+3\vec{j}-2\vec{k}\right)+n_3\left(2\vec{i}-4\vec{j}+x\vec{k}\right)=0\vec{i}+0\vec{j}+0\vec{k}\\ \small \left(2n_1+4n_2+2n_3\right)\vec{i}+\left(-3n_1+3n_2-4n_3\right)\vec{j}+\left(n_1-2n_2+xn_3\right)\vec{k}=0\vec{i}+0\vec{j}+0\vec{k}\] This gives you the system \[\begin{cases}\;\;\;2n_1+4n_2+2n_3=0\\ -3n_1+3n_2-4n_3=0\\ \;\;\;\;\;n_1-2n_2+xn_3=0\end{cases}\] Does that help?

OpenStudy (anonymous):

yes thanks but i dont know hwo to solve the simultaneous equation

OpenStudy (anonymous):

Do you know how to use matrices to describe the system? You are in linear algebra, correct?

OpenStudy (anonymous):

no this is high school. is it determinant ? oh i do know how to setup Ax=0

OpenStudy (anonymous):

The matrix equation is given by Ax=B, where B = 0 in this case. \[\left[\begin{matrix}2&4&2\\-3&3&-4\\1&-2&x\end{matrix}\right]\left[\begin{matrix}n_1\\n_2\\n_3\end{matrix}\right]=\left[\begin{matrix}0\\0\\0\end{matrix}\right]\]

OpenStudy (anonymous):

yeah need find det(a)=/=0 right

OpenStudy (anonymous):

wait det(a)=0 i mean

OpenStudy (anonymous):

Well, if the determinant of A is zero, then its inverse doesn't exist. To solve for each n, you need to have A's inverse.

OpenStudy (anonymous):

Do you know how to find the inverse of a 3x3 matrix?

OpenStudy (anonymous):

linear dependent if det=/=0 independet if =0

OpenStudy (anonymous):

no need calculator

OpenStudy (anonymous):

for "linear dependent if det=/=0 independet if =0" why does it work if you put the vectors as rows as well?

OpenStudy (anonymous):

Oh, disregard that, then. I wasn't aware of the "linear independence implies det=0" thing you mentioned. The way I wrote it helped me to find the determinant on paper, but I think it works in a similar way. You're going to have to know how to find the determinant by hand, though. A calculator won't let you put in x.

OpenStudy (anonymous):

i understand when you put as columns, if det=0, singular .: more than one solution and when det=/=0 invertible .: only trivial solution

OpenStudy (anonymous):

Right, I'm stuck in the mindset of solving equations in that way. Also, the "column" way of writing the vectors is just the transpose of the matrix I used, so their determinants are the same. Anyway, I get the determinant of A to be 18x - 28.

OpenStudy (anonymous):

can you explain why linear dependent if det=/=0 independet if =0 ??

OpenStudy (anonymous):

it works for row vectors as well because determinant the same ?

OpenStudy (anonymous):

whats transpose matrix

OpenStudy (anonymous):

The determinant of a square matrix is the same as that of its transpose. ?

OpenStudy (anonymous):

I just referred to this wiki page on linear independence: http://en.wikipedia.org/wiki/Linear_independence#Alternative_method_using_determinants The transpose of a matrix is a reflection of the matrix on itself, in a way. For example, the transpose of A would be denoted A^t, and \[A^T=\left[\begin{matrix}2&-3&1\\4&3&-2\\2&-4&x\end{matrix}\right]\] And yes, the determinant of A is the same as the determinant of A^T. But that's not important. What is important is that the system is linearly dependent if the determinant is zero.

OpenStudy (anonymous):

*NOT zero, I mean. Sorry.

OpenStudy (anonymous):

And another correction: My determinant should be 18x-26.

OpenStudy (anonymous):

okayokay wait so. for a system to be dependent, det=0? or =/=0

OpenStudy (anonymous):

are you sure "since the determinant is non-zero, the vectors (1, 1) and (−3, 2) are linearly independent." doesnt that say det=/=0, independent

OpenStudy (anonymous):

Agh, this wording is tripping me up. Should be For the system associated with matrix A to be linearly DEPENDENT, det(A) = 0. And that's final. I hope.

OpenStudy (anonymous):

yes thatswhat i thought

OpenStudy (anonymous):

can you explain why?

OpenStudy (anonymous):

I'm not sure how to explain it informally. Normally, I'd try to prove it for myself using what I've learned in linear algebra, and I don't think it'd be easy to express in a single post. For now, I would just accept it as fact for this particular problem, then investigate it later. The last thing you have to do is find the determinant, which is 18x-26. Do you know how to do that?

OpenStudy (anonymous):

nope. only learnt det for 2x2. teacher said to just use calculator for larger matrices

OpenStudy (anonymous):

but can you solve the 3 equations using gaussian elimination?

OpenStudy (anonymous):

need to do it by hand like that

OpenStudy (anonymous):

Yes, I think you could. You have to show that at least one of the n's isn't necessarily 0.

OpenStudy (anonymous):

i mean do you know hwo to do it becausei dont